A jet takes the same time to travel 2500 km with the wind as it does to travel 1900 km against the wind. If its speed relative to the air is 450 km/hr, what is the speed of the wind?
Answer by scott8148(6628)

t = d / r ___ the times are equal

2500 / (450 + w) = 1900 / (450 - w)

(25 * 450) - 25w = (19 * 450) + 19w

6 * 450 = 44w